# Device-independent certification of desirable properties with a confidence interval

### v1

@@ Abstract
In the development of quantum technologies, a reliable means for characterizing quantum devices, be it a measurement device, a state-preparation device, or a transformation device, is crucial. However, the conventional approach based on, for example, quantum state tomography or process tomography relies on assumptions that are often not necessarily justifiable in a realistic experimental setting. While the device-independent approach to this problem gets around the shortcomings above by making only minimal, justifiable assumptions, most of the theoretical proposals to date only work in the idealized setting where independent and identically distributed (i.i.d.) trials are assumed. Here, we provide a versatile solution for rigorous device-independent certification that does not rely on the i.i.d. assumption. Specifically, we describe how the prediction-based-ratio (PBR) protocol and martingale-based protocol developed for hypothesis testing can be applied in the present context to achieve a device-independent certification of desirable properties with confidence interval. To illustrate the versatility of these methods, we demonstrate how we can use them to certify—with finite data—the underlying negativity, Hilbert space dimension, entanglement depth, and fidelity to some target pure state. In particular, we give examples showing how the amount of certifiable negativity and fidelity scales with the number of trials. Our results also show that, while the martingale-based protocol is more straightforward to implement, its performance depends strongly on the choice of the Bell function. Intriguingly, a Bell function useful for self-testing does not necessarily give the optimal confidence-gain rate for certifying the fidelity to the corresponding target state.
@@ I Introduction
The proper analysis of quantum experiments is an indispensable part in the development of quantum technologies. However, it is not trivial to reliably characterize a quantum setup, which may include measurement devices and state-preparation devices. Moreover, imperfections in the experimental setup can easily result in a mismatch [1, 2, 3] between the characterization tools developed for an idealized situation and an actual experimental situation. However, we can circumvent this problem by the so-called “device-independent approach” [4, 5]. In quantum information, the term “device-independent” (DI) was first coined [6] in the task of quantum key distribution [7, 8, 9], even though the idea was already conceived independently, but implicitly in [10].
@@
In a nutshell, the DI approach is a framework for analyzing physical systems without relying on any assumption about the degrees of freedom measured. Its basis is Bell-nonlocality [11, 5], which shows that no local-hidden-variable theory (LHV) can reproduce all quantum predictions, even though no assumption is made about the details of such a theory. For example, it is known that the violation of Bell inequalities [11] obtained by locally measuring a shared state implies [12] shared entanglement [13], which is a powerful resource in many quantum information processing tasks. More generally, many other desirable properties of the underlying state [14, 15, 16, 17, 18, 19, 20, 21, 22], measurements [21, 22, 23, 24, 25, 26, 27], and channel [20, 28, 29] may be derived directly from the observation of a Bell-inequality-violating correlation between measurement outcomes.
@@
However, due to statistical fluctuations, even when the experimental trials are independent and identically distributed (i.i.d), relative frequencies of measurement outcomes obtained from a Bell experiment do not faithfully represent the underlying distribution. In particular, such raw distributions estimated from the experimental results typically [30, 31, 32] lead to a violation of the nonsignaling conditions [33, 34], which is a prerequisite for the analysis in [15, 16, 17, 18, 20, 21, 22, 23, 24, 25, 26, 28, 27, 29]. In other words, statistical fluctuations render the many theoretical tools developed for such a purpose inapplicable. To address this issue, some ad hoc methods [30, 31, 32] have been proposed to regularize the relative frequencies obtained to ensure that the resulting distribution satisfy the nonsignaling conditions. In [35], a more in-depth discussion was provided and two better-motivated regularization methods were proposed.
@@
While these more recent attempts do provide a point estimator that fits within the framework of the usual DI analysis, they are still problematic in two aspects. Firstly, they do not provide any confidence region associated with the estimate. However, any real experiment only involves a finite number of experimental trials. Therefore a useful analysis should provide not only an estimate but also an indication of the reliability of such an estimate. In many of the Bell experiments reported [36, 37, 38, 39], this is achieved by reporting the standard deviations of Bell violations. However, for finite, especially a relatively small number of trials, the central limit theorem is not warranted, so the usual interpretation of standard deviations may become dubious. Secondly, these usual approaches and those that provide a DI point estimator [30, 31, 32, 35] implicitly assumes that the experimental trials are i.i.d, and hence free of the memory effect [40, 41] (see more discussions in [5, 42, 43, 44]). Of course, in a realistic experimental setting, the i.i.d assumption may be difficult to justify.
@@
To this end, we structure the rest of this paper as follows. In Section II, we explain the basic concepts relevant to the understanding of DI certification in the ideal setting. After that, we introduce in Section III our adapted statistical tools for performing a rigorous device-independent certification. Results obtained from these tools are then presented in Section IV. Finally, we give our concluding remarks and future directions in Section V.
@@ II.1 Correlations and Bell inequalities
where the Bell coefficients $\beta^{ab}_{xy}\in\mathbb{R}$, $\vec{\beta}:=\{\beta^{ab}_{xy}\}$, $B_{\mathcal{L}}(\{\vec{\beta}\})$ is the so-called local (upper) bound, and we use $\mathcal{L}$ to signify that the inequality holds under the assumption that $\vec{P}$ is compatible with a LHV theory. Explicitly, the nature of such a theory demands that $\vec{P}$ is factorizable in the form of [11, 5]
$P(ab|xy)\overset{\mathcal{L}}{=}\sum_{\lambda}q_{\lambda}P(a|x\lambda)P(b|y\lambda)$ (2)
where $q_{\lambda}\geq 0$ for all $\lambda$, $\sum_{\lambda}q_{\lambda}=1$, and $P(a|x\lambda),P(b|y\lambda)\in[0,1]$ are local response functions.
In an actual Bell test, the measurement settings ought to be chosen randomly according to some predetermined distribution $P_{xy}$. To manifest this fact, one may write Eq. 1 using the unconditional joint distribution $P(a,b,x,y)=P(ab|xy)P_{xy}$. In turn, we can then write a Bell inequality as a bound on the expectation value of a Bell function $I(v)$, defined in terms of $\vec{\beta}$ and $P_{xy}$, i.e.,
$\left\langle I(v)\right\rangle:=\left\langle\tfrac{\beta^{ab}_{xy}}{P_{xy}}\right\rangle\overset{\mathcal{L}}{\leq}B_{\mathcal{L}}(\vec{\beta})$ (3)
@@
where $\{M^{(A)}_{a|x}\}$ and $\{M^{(B)}_{b|y}\}$ are, respectively, the local positive-operator-valued measure describing Alice and Bob’s local measurements. For the benefits of subsequent discussions, it is also worth noting that both LHV and quantum correlations satisfy the nonsignaling conditions [33, 34]:
$\begin{split}\sum_{a}P(ab|xy)=\sum_{a}P(ab|xy^{\prime})\quad\forall\,\,y,y^{\prime},\\ \sum_{b}P(ab|xy)=\sum_{b}P(ab|x^{\prime}y)\quad\forall\,\,x,x^{\prime}.\end{split}$ (7)
@@ II.2.1 Negativity and dimension
Also worth noting is that if $\rho$ acts on $\mathbb{C}^{d_{A}}\otimes\mathbb{C}^{d_{A}}$ with $d=\min\{d_{A},d_{B}\}$, then the maximal possible negativity $\mathcal{N}(\rho)$ is upper bounded by $\mathcal{N}^{d}_{\text{max}}:=\frac{d-1}{2}$. Consequently, the observation of a large enough negativity also provides a nontrivial lower bound on the local Hilbert space dimension of the underlying system. More precisely, if the lower bound of $\mathcal{N}(\rho)$ obtained from Section II.2.1 exceeds $\mathcal{N}^{d}_{\text{max}}$, one immediately deduces that $\rho$ must act on a local Hilbert space of dimension $\geq d+1$, thereby giving a dimension witness [14].
@@
$I_{\text{\tiny CGLMP3}}:\beta^{ab}_{xy}=$ $(-1)^{x(y-1)}[\delta^{(2)}_{a-b}-(1-\delta_{x}\delta_{y-1})\delta^{(3)}_{b-a-1}]$
$-\delta_{x}\delta_{y-1}\delta^{(3)}_{b-a+1}\quad\text{and}\quad B_{\mathcal{L}}(\vec{\beta})=2,$ (11)
@@ II.2.2 Entanglement depth
In a many-body system, entanglement can come in various forms or structures [72]. In particular, an $n$-partite quantum state that is not fully separable is not necessarily genuinely $n$-partite entangled either. To witness the latter, one could rely on the demonstration of so-called genuine multipartite nonlocality [73]. However, as remarked in [15], it is possible to witness so-called genuine multipartite entanglement without relying on this strong form of multipartite nonlocality. In fact, using the semidefinite program introduced in [16], one can even systematically construct DI witnesses of this kind, starting from a given multipartite Bell function, say $\vec{\beta}=\{\beta^{abc}_{xyz}\}$. Later, it was further shown in [18] (see also [74]) that the extent to which a multipartite Bell inequality is violated can be used to witness (lower-bound) the underlying entanglement depth [75, 76], i.e., the extent to which a many-body entanglement is needed to prepare the underlying multipartite state.
For illustration, consider the expectation value of the Mermin Bell function [77] $I_{\text{Mermin}}(v)$ with $v=(a,b,c,x,y,z)$:
$\mathcal{S}_{\text{\tiny Mermin}}=\left\langle I_{\text{Mermin}}(v)\right\rangle=\left\langle\frac{\beta^{abc}_{xyz}}{P_{xyz}}\right\rangle=\sum_{x,y,z}{}^{{}^{\prime}}(-1)^{xyz}E_{xyz}$ (12)
@@ II.2.3 State fidelity
The strongest form of device certification one can hope for within a DI paradigm is called self-testing [78], first proposed in [10]. The key observation behind this feat is that the quantum strategy compatible with certain extremal quantum correlations $\vec{P}_{\mathcal{Q}}$ is essentially unique. Hence, with the observation of $\vec{P}_{\mathcal{Q}}$ in a Bell test, we can conclude, unambiguously that some degree of freedom (DOF) of the measured system must match a specific target state |$\ket{\psi}$. Often, one can also self-test the underlying measurements alongside the state (see, however, [79, 80] for examples of some exceptions).
@@
$M^{(A)}_{a|x}=\frac{\mathbb{1}+(-1)^{a}A_{x}}{2},\quad M^{(B)}_{b|y}=\frac{\mathbb{1}+(-1)^{a}B_{y}}{2}.$ (16)
@@
|$\ket{\psi(\theta)}=\cos\theta\ket{00}+\sin\theta\ket{11},\quad\theta\in(0,\frac{\pi}{4})$ (17a)
it suffices [71] to consider $A_{x}$ of Eq. 15c but generalize $B_{y}$ to:
@@
Interestingly, the resulting correlations also self-test [85, 86] the corresponding quantum strategy of Section II.2.3 and maximally violate the family of tilted CHSH Bell inequalities for $\alpha=2\sqrt{\frac{\cos^{2}2\theta}{1+\sin^{2}2\theta}}$:
@@
In practice, however, due to various imperfections, one can, at best attains a correlation close to the ideal one $\vec{P}_{\mathcal{Q}}$. In other words, in a realistic experimental setting, one can only hope to lower bound the similarity of the measured state $\rho$ with respect to the target state |$\ket{\psi}$ via a fidelity measure. To this end, a powerful numerical technique known as the SWAP method has been introduced in [87] (see also [86]) for exactly this purpose. More precisely, for any observed quantum correlation $\vec{P}$, the method allows one to lower bound the fidelity:
@@
extracted from the underlying quantum state $\rho$ via some local extraction map $\Phi$, which is a function of the actual POVM elements. Consequently, $\mathcal{F}$ is a function of the entries of the moment matrix $\chi_{\ell}[\rho]$, discussed below Section II.2.1. For details of the method, we refer the readers to [86].
@@ II.3 Some general remarks
At this point, it is worth noting that for all the three properties $\mathcal{P}$ discussed above—negativity (and hence dimension), entanglement depth, and reference-state fidelity—their DI certification can be achieved via the characterization of some convex set $\mathcal{C}_{\mathcal{P}}$ in the space of correlation vectors $\{\vec{P}\}$. More precisely, for negativity, by turning the objective function of Section II.2.1 into the constraint [16]
@@
we obtain an SDP that characterizes of the set of correlations attainable by quantum states having a negativity upper-bounded by $\mathcal{N}_{0}$. On the other hand, if we drop the constraint in Section II.2.1 but imposes additional positive-partial-transposition constraints, then we get an SDP characterization of the set $\mathcal{C}_{\mathcal{P}}$ having a bounded amount of entanglement depth [18] [see constraints of Section IV.2 below]. Likewise, by demanding $\bra{\psi}\rho_{\text{\tiny SWAP}}\ket{\psi}\leq\mathcal{F}_{0}$ together with Eq. 9c, we obtain an SDP characterization of the set $\mathcal{C}_{\mathcal{P}}$ associated with a swapped state [86] that has a fidelity to |$\ket{\psi}$ that is upper bounded by $\mathcal{F}_{0}$. In fact, an SDP characterization can also be obtained for a number of other properties, including genuine negativity [16], steering robustness [21], entanglement robustness [22], (measurement) incompatibility robustness [21, 27], and so on.

### v2

@@ Abstract
In the development of quantum technologies, a reliable means for characterizing quantum devices, be it a measurement device, a state-preparation device, or a transformation device, is crucial. However, the conventional approach based on, for example, quantum state tomography or process tomography relies on assumptions that are often not necessarily justifiable in a realistic experimental setting. While the device-independent approach to this problem bypasses the shortcomings above by making only minimal, justifiable assumptions, most of the theoretical proposals to date only work in the idealized setting where independent and identically distributed (i.i.d.) trials are assumed. Here, we provide a versatile solution for rigorous device-independent certification that does not rely on the i.i.d. assumption. Specifically, we describe how the prediction-based-ratio (PBR) protocol and martingale-based protocol developed for hypothesis testing can be applied in the present context to achieve a device-independent certification of desirable properties with confidence interval. To illustrate the versatility of these methods, we demonstrate how we can use them to certify—with finite data—the underlying negativity, Hilbert space dimension, entanglement depth, and fidelity to some target pure state. In particular, we give examples showing how the amount of certifiable negativity and fidelity scales with the number of trials, and how many experimental trials one needs to certify a qutrit state space, or the presence of genuine tripartite entanglement. Overall, we have found that the PBR protocol and the martingale-based protocol often offer similar performance, even though the former does have to presuppose any witness (Bell-like inequality). In contrast, our findings also show that the performance of the martingale-based protocol may be severely affected by one’s choice of the Bell-like inequality. Intriguingly, a Bell function useful for self-testing does not necessarily give the optimal confidence-gain rate for certifying the fidelity to the corresponding target state.
@@ I Introduction
The proper analysis of quantum experiments is an indispensable part in the development of quantum technologies. However, it is not trivial to reliably characterize a quantum setup, which may include, e.g., measurement devices and state-preparation devices. Moreover, imperfections in the experimental setup can easily result in a mismatch [1, 2, 3] between the characterization tools developed for an idealized situation and an actual experimental situation. However, we can circumvent this problem by the so-called “device-independent approach” [4, 5]. In quantum information, the term “device-independent” (DI) was first coined [6] in the task of quantum key distribution [7, 8, 9], even though the idea was already conceived independently, but implicitly in [10, 11].
@@
In a nutshell, the DI approach is a framework for analyzing physical systems without relying on any assumption about the degrees of freedom measured. Its basis is Bell-nonlocality [12, 5], which shows that no local-hidden-variable theory (LHV) can reproduce all quantum predictions, even though no further assumption is made about the details of such a theory. For example, it is known that the violation of Bell inequalities [12] obtained by locally measuring a shared state implies [13] shared entanglement [14], which is a powerful resource in many quantum information processing tasks. More generally, many other desirable properties of the underlying state [15, 16, 17, 18, 19, 20, 21, 22, 23], measurements [22, 23, 24, 25, 26, 27, 28], and channel [21, 29, 30] may be derived directly from the observation of a Bell-inequality-violating correlation between measurement outcomes. Recently, the DI approach has been also been incorporated into the security analysis of quantum secure direct communication, see, e.g., [31] and references therein.
@@
However, due to statistical fluctuations, even when the experimental trials are independent and identically distributed (i.i.d), relative frequencies of the measurement outcomes obtained from a Bell experiment do not faithfully represent the underlying distribution. In particular, such raw distributions estimated from the experimental results typically [32, 33, 34] lead to a violation of the nonsignaling conditions [35, 36], which is a prerequisite for the analysis in [16, 17, 18, 19, 21, 22, 23, 24, 25, 20, 26, 27, 29, 28, 30]. In other words, statistical fluctuations render the many theoretical tools developed for such a purpose inapplicable. To address this issue, some ad hoc methods [32, 33, 34] have been proposed to regularize the relative frequencies obtained to ensure that the resulting distribution satisfy the nonsignaling conditions. In [37], a more in-depth discussion was provided and two better-motivated regularization methods were proposed.
@@
While these more recent attempts do provide a point estimator that fits within the framework of the usual DI analysis, they are still problematic in two aspects. Firstly, they do not provide any confidence region associated with the estimate. However, any real experiment necessarily involves only a finite number of experimental trials. Therefore a useful analysis should provide not only an estimate but also an indication of the reliability of such an estimate. In many of the Bell experiments reported [38, 39, 40, 41], this is achieved by reporting the standard deviations of Bell violations. However, for finite, especially a relatively small number of trials, the central limit theorem is not warranted, so the usual interpretation of standard deviations may become dubious. Secondly, these usual approaches and those that provide a DI point estimator [32, 33, 34, 37] implicitly assumes that the experimental trials are i.i.d, and hence free of the memory effect [42, 43] (see more discussions in [5, 44, 45, 46]). Again, in a realistic experimental setting, the i.i.d assumption may be difficult to justify.
@@
To this end, we structure the rest of this paper as follows. In Section II, we explain the basic concepts relevant to the understanding of DI certification in the ideal setting. After that, we introduce in Section III our adapted statistical tools for performing a rigorous device-independent certification. Results obtained from these tools are then presented in Section IV. Finally, we give some concluding remarks and future directions in Section V.
@@ II.1 Correlations and Bell inequalities
where the Bell coefficients $\beta^{ab}_{xy}\in\mathbb{R}$, $\vec{\beta}:=\{\beta^{ab}_{xy}\}$, and $B_{\mathcal{L}}(\{\vec{\beta}\})$ is the so-called local (upper) bound. Here, we use $\mathcal{L}$ to signify that the inequality holds under the assumption that $\vec{P}$ is compatible with an LHV theory. Explicitly, the nature of such a theory demands that $\vec{P}$ is factorizable in the form of [12, 5]
$P(ab|xy)\overset{\mathcal{L}}{=}\sum_{\lambda}q_{\lambda}P_{A}(a|x\lambda)P_{B}(b|y\lambda)$ (2)
where $q_{\lambda}\geq 0$ for all $\lambda$, $\sum_{\lambda}q_{\lambda}=1$, and $P_{A}(a|x\lambda),P_{B}(b|y\lambda)\in[0,1]$ are local response functions.
In an actual Bell test, the measurement settings ought to be chosen randomly according to some predetermined distribution $P_{xy}$. To manifest this fact, one may write Eq. 1 using the unconditional joint distribution $P(abxy)=P(ab|xy)P_{xy}$ such that $P_{xy}=\sum_{a,b}P(abxy)$. In turn, we can then write a Bell inequality as a bound on the expectation value of a Bell function $I(v)$, defined in terms of $\vec{\beta}$ and $P_{xy}$, i.e.,
$\left\langle I(v)\right\rangle:=\tfrac{\left\langle\beta^{ab}_{xy}\right\rangle}{P_{xy}}\overset{\mathcal{L}}{\leq}B_{\mathcal{L}}(\vec{\beta})$ (3)
@@
where $\{M^{(A)}_{a|x}\}$ and $\{M^{(B)}_{b|y}\}$ are, respectively, the local positive-operator-valued measure (POVM) describing Alice and Bob’s local measurements. For the benefits of subsequent discussions, it is also worth noting that both LHV and quantum correlations satisfy the nonsignaling conditions [35, 36]:
$\begin{split}\sum_{a}P(ab|xy)=\sum_{a}P(ab|x^{\prime}y)\quad\forall\,\,x,x^{\prime},\\ \sum_{b}P(ab|xy)=\sum_{b}P(ab|xy^{\prime})\quad\forall\,\,y,y^{\prime}.\end{split}$ (7)
@@ II.2.1 Negativity and dimension
Also worth noting is that if $\rho$ acts on $\mathbb{C}^{d_{A}}\otimes\mathbb{C}^{d_{B}}$ with $d=\min\{d_{A},d_{B}\}$, then the maximal possible negativity $\mathcal{N}(\rho)$ is upper bounded by $\mathcal{N}^{d}_{\text{max}}:=\frac{d-1}{2}$. Consequently, the observation of a large enough negativity also provides a nontrivial lower bound on the local Hilbert space dimension of the underlying system. More precisely, if the lower bound on $\mathcal{N}(\rho)$ obtained from Eq. 9 exceeds $\mathcal{N}^{d}_{\text{max}}$, one immediately deduces that $\rho$ must act on a local Hilbert space of dimension $\geq d+1$, thereby giving a dimension witness [15].
@@
$I_{\text{\tiny CGLMP3}}:\beta^{ab}_{xy}=$ $(-1)^{x(y-1)}\{\delta^{(2)}_{a-b}-[1-\delta^{(2)}_{x}\delta^{(2)}_{y-1}]\delta^{(3)}_{b-a-1}\}$
$-\delta^{(2)}_{x}\delta^{(2)}_{y-1}\delta^{(3)}_{b-a+1}\quad\text{and}\quad B_{\mathcal{L}}(\vec{\beta})=2,$ (11)
@@ II.2.2 Entanglement depth
In a many-body system, entanglement can come in various forms or structures [74]. In particular, an $n$-partite quantum state that is not fully separable is not necessarily genuinely $n$-partite entangled either. To witness the latter, one could rely on the demonstration of so-called genuine multipartite nonlocality [75]. However, as remarked in [16], it is possible to witness genuine multipartite entanglement without relying on this strong form of multipartite nonlocality. In fact, using the SDP introduced in [17], one can even systematically construct DI witnesses of this kind, starting from a given multipartite Bell function, say $\vec{\beta}=\{\beta^{abc}_{xyz}\}$. Later, it was further shown in [19] (see also [76]) that the extent to which a multipartite Bell inequality is violated can be used to witness (lower-bound) the underlying entanglement depth [77, 78], i.e., the extent to which a many-body entanglement is needed to prepare the given multipartite state.
For illustration, consider the expectation value of the Mermin Bell function [79] $I_{\text{\tiny Mermin}}(v)$ with $v=(a,b,c,x,y,z)$:
$\mathcal{S}_{\text{\tiny Mermin}}=\left\langle I_{\text{\tiny Mermin}}(v)\right\rangle=\frac{\left\langle\beta^{abc}_{xyz}\right\rangle}{P_{xyz}}=\sum_{x,y,z}{}^{{}^{\prime}}(-1)^{xyz}E_{xyz}$ (12)
@@ II.2.3 State fidelity
The strongest form of device certification one can hope for within a DI paradigm is called self-testing [80], first proposed in [10]. The key observation behind this feat is that the quantum strategy compatible with certain extremal quantum correlations $\vec{P}_{\mathcal{Q}}$ is essentially unique. Hence, with the observation of $\vec{P}_{\mathcal{Q}}$ in a Bell test, we can conclude, unambiguously that some degree of freedom (DOF) of the measured system must match a specific target state |$\ket{\psi}$. Often, one can also self-test the underlying measurements alongside the state (see, however, [81, 82] for some examples of exceptions).
@@
$M^{(A)}_{a|x}=\frac{\mathbb{1}+(-1)^{a}A_{x}}{2},\quad M^{(B)}_{b|y}=\frac{\mathbb{1}+(-1)^{b}B_{y}}{2}.$ (16)
@@
|$\ket{\psi(\theta)}=\cos\theta\ket{00}+\sin\theta\ket{11},\quad\theta\in\left(0,\frac{\pi}{4}\right]$ (17a)
it suffices [73] to consider $A_{x}$ of Eq. 15c but generalize $B_{y}$ to [87]:
@@
Interestingly, the resulting correlation also self-tests [88, 87] the corresponding quantum strategy of Eq. 17 and maximally violate the family of tilted CHSH Bell inequalities for $\alpha=2\sqrt{\frac{\cos^{2}2\theta}{1+\sin^{2}2\theta}}$:
@@
In practice, however, due to various imperfections, one can, at best attain a correlation close to the ideal one $\vec{P}_{\mathcal{Q}}$. In other words, in a realistic experimental setting, one can only hope to lower bound the similarity of the measured state $\rho$ with respect to the target state |$\ket{\psi}$ via a fidelity measure. To this end, a powerful numerical technique known as the SWAP method has been introduced in [89] (see also [87]) for exactly this purpose. More precisely, for any observed quantum correlation $\vec{P}$, the method allows one to lower bound the fidelity:
@@
extracted from the underlying quantum state $\rho$ via some local extraction map $\Phi$, which is a function of the actual POVM elements. Consequently, $\mathcal{F}$ is a function of the entries of the moment matrix $\chi_{\ell}[\rho]$, discussed below Eq. 9. For details of the method, we refer the readers to [87].
@@ II.3 Some general remarks
At this point, it is worth noting that for all the three properties $\mathcal{P}$ discussed above—negativity (and hence dimension), entanglement depth, and reference-state fidelity—their DI certification can be achieved via the characterization of some convex set $\mathcal{C}_{\mathcal{P}}$ in the space of correlation vectors $\{\vec{P}\}$. More precisely, for negativity, by turning the objective function of Eq. 9 into the constraint [17]
@@
we obtain an SDP that characterizes of the set of correlations attainable by quantum states having a negativity upper-bounded by $\mathcal{N}_{0}$. Then, Eq. 10 can be understood as a separating hyperplane relevant for witnessing a negativity larger than $\mathcal{N}_{0}$.
On the other hand, if we drop the constraint in Eq. 9 but imposes additional positive-partial-transposition constraints, then we get an SDP characterization of the set $\mathcal{C}_{\mathcal{P}}$ having a bounded amount of entanglement depth [19] [see constraints of Eq. 44 below]. In this case, the first two inequalities of Eq. 14 serve as the corresponding witness for entanglement depth. Finally, by demanding $\bra{\psi}\rho_{\text{\tiny SWAP}}\ket{\psi}\leq\mathcal{F}_{0}$ together with Eq. 9c, we obtain an SDP characterization of the set $\mathcal{C}_{\mathcal{P}}$ associated with a swapped state [87] with a |$\ket{\psi}$-fidelity upper bounded by $\mathcal{F}_{0}$. In fact, an SDP characterization can also be obtained for a number of other properties, including genuine negativity [17], steering robustness [22], entanglement robustness [23], (measurement) incompatibility robustness [22, 28], and so on.